OpenStudy (anonymous):

What is the geometric mean between 12 and 3?

OpenStudy (anonymous):

6

OpenStudy (saifoo.khan):

add and divide.

OpenStudy (anonymous):

sqrt of 12 * 3

OpenStudy (anonymous):

heck no!

OpenStudy (anonymous):

@saifoo its goemetric mean

OpenStudy (anonymous):

oh sorry not you sivanand. make saifoo take a break. he is trying to get the land speed record

OpenStudy (anonymous):

\[\sqrt{12\times 3}=\sqrt{36}=6\]

OpenStudy (saifoo.khan):

Nothing happend, lol.